# Leave-One-Out Neighborhood Smoothing for Graphons: Berry-Esseen Bounds, Confidence Intervals, and Honest Tuning

### v1

@@ 3.3 Row‑wise risk
Under Assumptions 2.1 and 2.2 and with $h$ chosen as in the main text, there exist constants $C,c>0$ such that, with probability at least $1-n^{-c}$ and uniformly over $1\leq i\leq n$,
$R_{i}:=\frac{1}{n}\sum_{j=1}^{n}(\tilde{P}_{ij}-P_{ij})^{2}\leq C\sqrt{\frac{\log n}{n}}.$
Proof sketch.
Decompose $\tilde{P}_{ij}-P_{ij}=U_{ij}+B_{ij}$. By Lemma 3.1 and a union bound, $\max_{j}|U_{ij}|\lesssim\sqrt{\log n/n}$ with high probability, which implies $\frac{1}{n}\sum_{j}U_{ij}^{2}\lesssim\log n/n$. Lemma 3.2 gives $\frac{1}{n}\sum_{j}B_{ij}^{2}\lesssim(\log n/n)^{1/2}$. Comparing these two contributions yields the stated rate. Full details are deferred to Appendix D. ∎
Corollary 3.4 (Minimax rate match).
Under the conditions of Theorem 3.3, the LOO neighborhood smoother achieves the same row-wise mean squared error rate as classical neighborhood smoothing:
$\frac{1}{n}\sum_{j=1}^{n}(\tilde{P}_{ij}-P_{ij})^{2}\lesssim\sqrt{\frac{\log n}{n}}.$
In particular, up to logarithmic factors the LOO estimator attains the minimax rate for piecewise Lipschitz graphons studied in Gao et al. (2015); Zhang et al. (2017).
@@ 3.4 Berry-Esseen normal approximation
To obtain normal‑approximation intervals we apply a conditional Berry-Esseen bound to the sum defining $U_{ij}$. Let
$\sigma_{ij}^{2}:=\sum_{k\in\mathcal{N}_{i}^{(-j)}}\mathrm{Var}(A_{kj}-P_{kj})=\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj}),$
so that $V_{ij}=h^{-2}\sigma_{ij}^{2}$. Under Assumption 2.2 and the neighborhood construction one shows $\sigma_{ij}^{2}\asymp h$ uniformly over $(i,j)$ with high probability; hence the conditional variance of $U_{ij}$ is of order $h^{-1}$ and does not degenerate.
Theorem 3.5 (Berry-Esseen bound).
Suppose Assumptions 2.1 and 2.2 hold and $h$ is chosen as in the main text. Let $U_{ij}$ and $V_{ij}$ be as in (3.1)-(3.3). Then, with probability tending to one and uniformly over $1\leq i,j\leq n$,
$\sup_{x\in\mathbb{R}}\Big|\mathbb{P}\!\Big(\frac{U_{ij}}{\sqrt{V_{ij}}}\leq x\ \Big|\ \mathcal{N}_{i}^{(-j)},P\Big)-\Phi(x)\Big|\leq Ch^{-1/2}.$ (3.5)
In particular, with the canonical choice $h\asymp\sqrt{n\log n}$ the Berry-Esseen error decays as $O(h^{-1/2})=O\big((n\log n)^{-1/4}\big)$, which is comparable to the bias scale up to logarithmic factors.
Sketch of proof.
Conditional on $(\mathcal{N}_{i}^{(-j)},P)$ let $X_{k}=A_{kj}-P_{kj}$ for $k\in\mathcal{N}_{i}^{(-j)}$. These are independent, mean zero and bounded by one. Define $\rho_{3}=\sum_{k}\mathbb{E}[|X_{k}|^{3}\mid\mathcal{N}_{i}^{(-j)},P]$. Since $|X_{k}|\leq 1$ we have $\rho_{3}\leq\sigma_{ij}^{2}$. The classical Berry-Esseen theorem for independent, non‑identically distributed summands (see Petrov (1975); Vershynin (2018)) gives a bound of order $\rho_{3}/\sigma_{ij}^{3}\lesssim\sigma_{ij}^{-1}$. Using $\sigma_{ij}^{2}\asymp h$ and rescaling by $h^{-1}$ yields (3.5). Full details are given in Appendix A. ∎
If $h\asymp\sqrt{n\log n}$ the Berry-Esseen error in (3.5) is $O\big((n\log n)^{-1/4}\big)$, which is comparable to the bias scale in Lemma 3.2 up to logarithmic factors.

### v3

@@ 3.1 Neighborhood bias
Then there exist constants $C,c>0$ such that with probability at least $1-n^{-c}$, uniformly over $i$,
$\frac{1}{n}\sum_{j\neq i}\big(\tilde{P}^{\mathrm{ZLZ}}_{ij}-P_{ij}\big)^{2}\;\leq\;C\Big(\frac{\log n}{n}\Big)^{1/2}.$
Proof.
This is the row-wise conclusion proved for $\tilde{P}_{ij}$ in Zhang et al. (2017), using their supplementary decomposition into $J_{1}(i,j)$ and $J_{2}(i,j)$ (equation 1 in the supplementary material of Zhang et al. (2017)) and the two-hop concentration input. ∎
Corollary 3.4 (Leave-one-out ZLZ row-wise risk + $(U_{ij},B_{ij})$ decomposition).
Fix $j\in[n]$. Let $A^{(-j)}$ be the adjacency matrix with row/column $j$ deleted, and let $\widetilde{P}^{(-j)}$ denote the ZLZ neighborhood-smoothing estimator computed from $A^{(-j)}$ (with neighborhoods $N_{i}^{(-j)}$ defined exactly as in ZLZ but using the two-hop matrix $(A^{(-j)})^{2}/(n-1)$). For $i\neq j$ and $k\in N_{i}^{(-j)}$, define
$U^{(-j)}_{ij}:=\frac{1}{|N_{i}^{(-j)}|}\sum_{k\in N_{i}^{(-j)}}(A_{kj}-P_{kj}),\qquad B^{(-j)}_{ij}:=\frac{1}{|N_{i}^{(-j)}|}\sum_{k\in N_{i}^{(-j)}}(P_{kj}-P_{ij}),$
so that
$\widetilde{P}^{(-j)}_{ij}-P_{ij}=U^{(-j)}_{ij}+B^{(-j)}_{ij}$.
Then for all $i\neq j$,
$\frac{1}{n-1}\sum_{\ell\neq j}\big(\widetilde{P}^{(-j)}_{i\ell}-P_{i\ell}\big)^{2}\;\leq\;2\,\frac{1}{n-1}\sum_{\ell\neq j}\big(U^{(-j)}_{i\ell}\big)^{2}\;+\;2\,\frac{1}{n-1}\sum_{\ell\neq j}\big(B^{(-j)}_{i\ell}\big)^{2},$
and, with probability at least $1-c_{1}n^{-c_{2}}$ (for the same type of constants as ZLZ), uniformly for all $i\neq j$,
$\frac{1}{n-1}\sum_{\ell\neq j}\big(\widetilde{P}^{(-j)}_{i\ell}-P_{i\ell}\big)^{2}\;\lesssim\;\Big(\frac{\log n}{n}\Big)^{1/2}.$
Moreover, identifying ZLZ’s terms by
$J_{1}^{(-j)}(i,\ell):=\big(U^{(-j)}_{i\ell}\big)^{2},\qquad J_{2}^{(-j)}(i,\ell):=\big(B^{(-j)}_{i\ell}\big)^{2},$
one has the ZLZ-style bounds (with (n1)(n-1) in place of $n$):
$\frac{1}{n-1}\sum_{\ell\neq j}J_{1}^{(-j)}(i,\ell)\;\lesssim\;\Big(\frac{\log n}{n}\Big)^{1/2},\qquad\frac{1}{n-1}\sum_{\ell\neq j}J_{2}^{(-j)}(i,\ell)\;\lesssim\;\Big(\frac{\log n}{n}\Big)^{1/2}.$
Proof.
First, (ZLZ algebra; defines $J_{1},J_{2}$ and matches $U,B$). Exactly as in ZLZ, we begin from the decomposition of the squared error into two contributions:
$\Big(\widetilde{P}^{(-j)}_{i\ell}-P_{i\ell}\Big)^{2}=\Bigg[\frac{1}{|N_{i}^{(-j)}|}\sum_{k\in N_{i}^{(-j)}}\Big\{(A_{k\ell}-P_{k\ell})+(P_{k\ell}-P_{i\ell})\Big\}\Bigg]^{2},$
and the inequality
$(\cdot)^{2}\leq 2J_{1}^{(-j)}(i,\ell)+2J_{2}^{(-j)}(i,\ell)$
is the same inequality as ZLZ equation (1), with the identifications $J_{1}^{(-j)}(i,\ell)=(U^{(-j)}_{i\ell})^{2}$ and $J_{2}^{(-j)}(i,\ell)=(B^{(-j)}_{i\ell})^{2}$. (See equation (1), supplementary material of Zhang et al. (2017) .)
Second, two-hop concentration under deletion preserves ZLZ inputs. ZLZ’s neighborhood construction and the control of $J_{2}$ rely on uniform control of two-hop quantities of the form $(A^{2}/n)_{ab}-(P^{2}/n)_{ab}$, established in ZLZ equations, (4)-(6) in their supplementary material. By Lemma 3.2, the same bound holds when replacing $A$ by $A^{(-j)}$ (and $n$ by n1n-1). Therefore, every subsequent ZLZ step that uses (4)-(6) remains valid for the graph with the deleted vertex.
Third, Bounding the $J_{1}$ row-average ZLZ does not bound each $J_{1}(i,\ell)$ separately; instead it controls the row-average $(n^{-1}\sum_{\ell}J_{1}(i,\ell))$ via the decomposition in equation (11), together with the bounds (12), (14), and the conclusion (15) in the ZLZ supplementary material. The same argument applies verbatim to $A^{(-j)}$ and $N_{i}^{(-j)}$, using equation 3.2 to justify the same concentration inputs. Hence
$\frac{1}{n-1}\sum_{\ell\neq j}J_{1}^{(-j)}(i,\ell)\;\lesssim\;\Big(\frac{\log n}{n}\Big)^{1/2},$
by the ZLZ chain (11)-(15).
Fourth, ZLZ bounds $(n^{-1}\sum_{\ell}J_{2}(i,\ell))$ using Cauchy-Schwarz and the “claim 2 of their Lemma 2” input, concluding in equation (16). The same reasoning applies here (again by Lemma 3.2 to justify the same concentration inputs), yielding the same order bound:
$\frac{1}{n-1}\sum_{\ell\neq j}J_{2}^{(-j)}(i,\ell)\;\lesssim\;\Big(\frac{\log n}{n}\Big)^{1/2}.$
This is exactly ZLZ equation (16) with the leave-one-out objects.
Finally, we have the same claimed row-wise risk bound, matching ZLZ’s stated goal (their inequality right after introducing $\widetilde{P}$, i.e., the display in the supplementary around the start of the proof of Theorem 1). ∎
Assumption 3.5 (Piecewise bi-Lipschitz smoothness).
The graphon $W:[0,1]^{2}\to[0,1]$ admits a finite vertical partition $0=b_{0}<b_{1}<\cdots<b_{K}=1$ with blocks $I_{m}:=(b_{m},b_{m+1})$ such that:
  1. 1.
    (Within-block Lipschitz in the first argument.) For all $m$ and all $u,u^{\prime}\in I_{m}$, $x\in[0,1]$,
    $|W(u,x)-W(u^{\prime},x)|\leq L|u-u^{\prime}|.$
  2. 2.
    (Across-block $L^{2}$ separation.) There exist $\eta,\delta>0$ such that for any $u\in I_{m}$ and $u^{\prime}\in I_{m^{\prime}}$ with $m\neq m^{\prime}$,
    $\lambda\big(\{x\in[0,1]:|W(u,x)-W(u^{\prime},x)|\geq\eta\}\big)\geq\delta.$
  3. 3.
    (Within-block $L^{2}$ identifiability.) For any $\varepsilon>0$ with $b_{m}+\varepsilon<b_{m+1}-\varepsilon$, there exists $\kappa(\varepsilon)>0$ such that for all $u,u^{\prime}\in[b_{m}+\varepsilon,b_{m+1}-\varepsilon]$,
    $\int_{0}^{1}(W(u,x)-W(u^{\prime},x))^{2}\,dx\;\geq\;\kappa(\varepsilon)\,|u-u^{\prime}|^{2}.$
Lemma 3.6 (Rowwise $L^{2}$ closeness implies latent closeness).
Assume Assumption 3.5. Fix $\varepsilon>0$ and define
$\mathcal{I}_{\varepsilon}=\bigcup_{m=0}^{K-1}[b_{m}+\varepsilon,b_{m+1}-\varepsilon].$
Let $U_{i},U_{k}\in\mathcal{I}_{\varepsilon}$ and set $P_{ab}=W(U_{a},U_{b})$. Then there exist constants $c_{0},c_{1}>0$ (depending only on the constants in Assumption 3.5 and on $\varepsilon$) such that, for all sufficiently large $n$,
$\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\;\leq\;c_{0}\Big(\frac{\log n}{n}\Big)^{1/2}\quad\Longrightarrow\quad\begin{cases}U_{i},U_{k}\text{ lie in the same block }I_{m},\\ |U_{i}-U_{k}|\leq c_{1}(\log n/n)^{1/4}.\end{cases}$
Proof.
Write $f_{i}(x)=W(U_{i},x)$ and $f_{k}(x)=W(U_{k},x)$.
We can easily see that if $U_{i}$ and $U_{k}$ lie in different blocks, then by Assumption 3.5(ii),
$\int_{0}^{1}(f_{i}-f_{k})^{2}\;\geq\;\eta^{2}\delta.$
For $c_{0}>0$ small enough and $n$ large, $c_{0}(\log n/n)^{1/2}<\eta^{2}\delta/2$, so the assumed upper bound on $\frac{1}{n}\sum_{\ell}(P_{i\ell}-P_{k\ell})^{2}$ (which concentrates around $\int_{0}^{1}(f_{i}-f_{k})^{2}$ conditional on $(U_{\ell})$) is incompatible with this. Hence $U_{i},U_{k}$ must lie in the same block $I_{m}$.
On the other hand, since $U_{i},U_{k}\in\mathcal{I}_{\varepsilon}$, we have $U_{i},U_{k}\in[b_{m}+\varepsilon,b_{m+1}-\varepsilon]$ for some $m$, and Assumption 3.5(iii) gives
$\int_{0}^{1}(f_{i}-f_{k})^{2}=\int_{0}^{1}(W(U_{i},x)-W(U_{k},x))^{2}\,dx\;\geq\;\kappa(\varepsilon)\,|U_{i}-U_{k}|^{2}.$
Combining this with the assumed upper bound yields
$\kappa(\varepsilon)\,|U_{i}-U_{k}|^{2}\;\leq\;c_{0}\Big(\frac{\log n}{n}\Big)^{1/2},$
so
$|U_{i}-U_{k}|\;\leq\;\sqrt{\frac{c_{0}}{\kappa(\varepsilon)}}\,\Big(\frac{\log n}{n}\Big)^{1/4}.$
Set $c_{1}:=\sqrt{c_{0}/\kappa(\varepsilon)}$. ∎
Lemma 3.7 (Bias bound for the LOO neighborhood smoother).
Assume Assumption 3.5. Let $h_{n}\asymp\sqrt{n\log n}$ and define $r_{n}=(\log n/n)^{1/4}$. For $i\neq j$, set
$B^{(-j)}_{ij}=\frac{1}{|\mathcal{N}_{i}^{(-j)}|}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(P_{kj}-P_{ij}),\qquad P_{ab}=W(U_{a},U_{b}).$
Suppose there exist constants $C_{2},c>0$ such that, with probability at least $1-n^{-c}$, the LOO neighborhoods satisfy
$\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\;\leq\;C_{2}\Big(\frac{\log n}{n}\Big)^{1/2}\qquad\text{for all }i\neq j\text{ and all }k\in\mathcal{N}_{i}^{(-j)}.$ (3.4)
Then for all $i\neq j$ with $U_{i},U_{j}\in\mathcal{I}_{\varepsilon}$,
$|B^{(-j)}_{ij}|\;\leq\;C\,r_{n},$
for some constant $C>0$ depending only on $(L,\eta,\delta,\kappa(\varepsilon),C_{2})$.
Proof.
Fix $i\neq j$ with $U_{i},U_{j}\in\mathcal{I}_{\varepsilon}$ and work on the event where (3.4) holds. Define
$\mathcal{B}_{i}=\Big\{k\in\mathcal{N}_{i}^{(-j)}:\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}>M\Big(\frac{\log n}{n}\Big)^{1/2}\Big\},$
for some constant $M>C_{2}$.
Averaging $(\star)$ over $k\in\mathcal{N}_{i}^{(-j)}$ gives
$\frac{1}{|\mathcal{N}_{i}^{(-j)}|}\sum_{k\in\mathcal{N}_{i}^{(-j)}}\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\;\leq\;C_{2}\Big(\frac{\log n}{n}\Big)^{1/2}.$
If $|\mathcal{B}_{i}|\geq 1$, then the left-hand side is at least
$\frac{|\mathcal{B}_{i}|}{|\mathcal{N}_{i}^{(-j)}|}M\Big(\frac{\log n}{n}\Big)^{1/2}\;\geq\;\frac{M}{|\mathcal{N}_{i}^{(-j)}|}\Big(\frac{\log n}{n}\Big)^{1/2},$
which is strictly larger than $C_{2}(\log n/n)^{1/2}$ for $M>C_{2}$ and $|\mathcal{N}_{i}^{(-j)}|\geq 1$, a contradiction. Hence $|\mathcal{B}_{i}|=0$, and for every $k\in\mathcal{N}_{i}^{(-j)}$,
$\frac{1}{n}\sum_{\ell=1}^{n}(P_{i\ell}-P_{k\ell})^{2}\;\leq\;M\Big(\frac{\log n}{n}\Big)^{1/2}.$
Applying Lemma 3.6 to each such $k$ yields that $U_{k}$ lies in the same block as $U_{i}$ and
$|U_{k}-U_{i}|\leq C^{\prime}r_{n}$
for some constant $C^{\prime}>0$. Since $U_{j}\in\mathcal{I}_{\varepsilon}$, the points $(U_{i},U_{j})$ and $(U_{k},U_{j})$ lie in the same Lipschitz region of $W$ in the first argument, so by Assumption 3.5(i),
$|P_{kj}-P_{ij}|=|W(U_{k},U_{j})-W(U_{i},U_{j})|\leq L|U_{k}-U_{i}|\leq LC^{\prime}r_{n}.$
Averaging over $k\in\mathcal{N}_{i}^{(-j)}$ gives
$|B^{(-j)}_{ij}|\;\leq\;Cr_{n}$
for some $C>0$, as claimed. ∎
Lemma 3.7 establishes a bias bound for the leave-one-out neighborhood smoother, showing that under the piecewise bi-Lipschitz smoothness assumption, the bias term $B_{ij}^{(-j)}$ is of order $r_{n}=(\log n/n)^{1/4}$ for nodes $i$ and $j$ whose latent positions lie in the interior of the same block. This result relies on the control of the average two-hop distance between node $i$ and its neighbors in the leave-one-out neighborhood, as well as on the Lipschitz continuity of the graphon within blocks.
3.2 Variance Bounds, Berry-Esseen Approximation, and Confidence Intervals
We introduce the conditional variance for $U_{ij}^{(-j)}$:
$V_{ij}:=\mathrm{Var}\!\big(U_{ij}^{(-j)}\mid\mathcal{N}_{i}^{(-j)},P\big)=\frac{1}{h^{2}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj}),$ (3.5)
so that $V_{ij}\leq 1/(4h)$. While $V_{ij}$ is analytically useful and measurable with respect to the $\sigma$-field generated by $(\mathcal{N}_{i}^{(-j)},P)$, it depends on the unknown latent probabilities $\{P_{kj}\}$.
To obtain tight, variance-adaptive control of $U_{ij}^{(-j)}$ using only observable quantities, we define the unbiased sample variance of the edges in the leave-one-out neighborhood:
$\hat{\sigma}_{ij}^{2}:=\frac{1}{h(h-1)}\sum_{\begin{subarray}{c}k,\ell\in\mathcal{N}_{i}^{(-j)}\\ k<\ell\end{subarray}}(A_{kj}-A_{\ell j})^{2}=\frac{h}{h-1}\,\tilde{P}_{ij}(1-\tilde{P}_{ij}),$ (3.6)
where $\tilde{P}_{ij}=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}A_{kj}$ is the empirical mean of the observed edges. Instead of relying on classical Bernstein bounds that require the true variance $V_{ij}$, we utilize the Empirical Bernstein inequality (Theorem 4 of Maurer and Pontil (2009)), which directly incorporates the sample variance $\hat{\sigma}_{ij}^{2}$. The key advantage of this approach is that it completely bypasses the need for asymptotic plug-iin justifications, providing a strictly finite-sample, variance-adaptive control of $U_{ij}^{(-j)}$ based entirely on the observable quantity $\hat{\sigma}_{ij}^{2}$.
Lemma 3.8 (Empirical Bernstein inequality).
For any $t>0$,
$\mathbb{P}\Bigg(\big|U_{ij}^{(-j)}\big|\geq\sqrt{\frac{2\hat{\sigma}_{ij}^{2}\,t}{h}}+\frac{7t}{3(h-1)}\ \Bigg|\ \mathcal{N}_{i}^{(-j)},P\Bigg)\leq 2e^{-t}.$
Sketch of proof.
Conditional on $(\mathcal{N}_{i}^{(-j)},P)$, the variables $\{A_{kj}:k\in\mathcal{N}_{i}^{(-j)}\}$ are independent and bounded in $[0,1]$. Their empirical mean is
$\tilde{P}_{ij}=\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}A_{kj},$
and its conditional expectation is
$\frac{1}{h}\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}.$
Thus $U_{ij}$ is exactly the deviation of the empirical mean from its conditional expectation. Applying Theorem 4 of Maurer and Pontil (2009) to the variables $X_{k}:=A_{kj}$, $k\in\mathcal{N}_{i}^{(-j)}$, and taking a union bound over both tails yields the stated inequality. ∎
Lemma 3.8 provides a strictly finite-sample, variance-adaptive control of $U_{ij}$ directly in terms of the empirical variance $\hat{\sigma}_{ij}^{2}$, thereby completely bypassing the need for asymptotic plug-in justifications.
@@ 3.3 Berry-Esseen normal approximation
We now establish a normal approximation for the variance component $U_{ij}^{(-j)}$ defined in (3.1). Recall
$U_{ij}^{(-j)}=\frac{1}{h_{n}}\sum_{k\in\mathcal{N}_{i}^{(-j)}}(A_{kj}-P_{kj}),$
and define the conditional variance
$\sigma_{ij}^{2}:=\sum_{k\in\mathcal{N}_{i}^{(-j)}}\mathrm{Var}(A_{kj}\mid P)=\sum_{k\in\mathcal{N}_{i}^{(-j)}}P_{kj}(1-P_{kj}),$
so that
$V_{ij}=\mathrm{Var}(U_{ij}^{(-j)}\mid\mathcal{N}_{i}^{(-j)},P)=\frac{1}{h_{n}^{2}}\sigma_{ij}^{2}.$
Under Assumption 2.2, there exists $\eta\in(0,1/2)$ such that
$\eta\leq P_{kj}\leq 1-\eta\qquad\text{for all }k,j.$
Hence
$\eta(1-\eta)\leq P_{kj}(1-P_{kj})\leq\frac{1}{4}.$
Since $|\mathcal{N}_{i}^{(-j)}|=h_{n}$, it follows that
$\eta(1-\eta)\,h_{n}\leq\sigma_{ij}^{2}\leq\frac{1}{4}\,h_{n}.$ (3.7)
In particular,
$\sigma_{ij}\geq\sqrt{\eta(1-\eta)}\,h_{n}^{1/2},\qquad\sigma_{ij}^{-1}\leq C\,h_{n}^{-1/2}$ (3.8)
for a constant $C>0$ independent of $(i,j,n)$. Consequently,
$V_{ij}=\frac{\sigma_{ij}^{2}}{h_{n}^{2}}\asymp h_{n}^{-1},$
and the conditional variance of $U_{ij}^{(-j)}$ does not degenerate.
Theorem 3.9 (Uniform Berry-Esseen bound).
Suppose Assumptions 2.1 and 2.2 hold and $h_{n}\asymp\sqrt{n\log n}$. Then there exists a universal constant $C>0$, independent of $i,j,n$, such that, uniformly over $1\leq i,j\leq n$,
$\sup_{x\in\mathbb{R}}\left|\mathbb{P}\!\left(\frac{U_{ij}^{(-j)}}{\sqrt{V_{ij}}}\leq x\ \middle|\ \mathcal{N}_{i}^{(-j)},P\right)-\Phi(x)\right|\leq Ch_{n}^{-1/2}.$ (3.9)
Proof.
Fix $(i,j)$ and condition on $(\mathcal{N}_{i}^{(-j)},P)$. Define
$X_{k}:=A_{kj}-P_{kj},\qquad k\in\mathcal{N}_{i}^{(-j)}.$
By Proposition 2.7, the variables $\{X_{k}\}$ are independent, mean zero and bounded in [1,1][-1,1], but not necessarily identically distributed.
Let
$S:=\sum_{k\in\mathcal{N}_{i}^{(-j)}}X_{k},\qquad\sigma_{ij}^{2}=\mathrm{Var}(S\mid\mathcal{N}_{i}^{(-j)},P).$
Define the third absolute moment sum
$\rho_{3}:=\sum_{k\in\mathcal{N}_{i}^{(-j)}}\mathbb{E}\!\left[|X_{k}|^{3}\mid\mathcal{N}_{i}^{(-j)},P\right].$
Since $|X_{k}|\leq 1$, we have
$\rho_{3}\leq\sum_{k}\mathbb{E}[X_{k}^{2}\mid\mathcal{N}_{i}^{(-j)},P]=\sigma_{ij}^{2}.$
The Berry-Esseen inequality for independent, not necessarily identically distributed summands (see, e.g., Feller (1971, Chapter XVI)) states that there exists a universal constant $C_{0}>0$ such that
$\sup_{x\in\mathbb{R}}\left|\mathbb{P}\!\left(\frac{S}{\sigma_{ij}}\leq x\ \middle|\ \mathcal{N}_{i}^{(-j)},P\right)-\Phi(x)\right|\leq C_{0}\frac{\rho_{3}}{\sigma_{ij}^{3}}.$
Using $\rho_{3}\leq\sigma_{ij}^{2}$ yields
$\sup_{x\in\mathbb{R}}\left|\mathbb{P}\!\left(\frac{S}{\sigma_{ij}}\leq x\ \middle|\ \mathcal{N}_{i}^{(-j)},P\right)-\Phi(x)\right|\leq C_{0}\sigma_{ij}^{-1}.$
By (3.8), $\sigma_{ij}^{-1}\leq Ch_{n}^{-1/2}$ and therefore it does not depend on $i$ and $j$. Finally, since
$U_{ij}^{(-j)}=\frac{1}{h_{n}}S,\qquad V_{ij}=\frac{1}{h_{n}^{2}}\sigma_{ij}^{2},$
we have the identity
$\frac{U_{ij}^{(-j)}}{\sqrt{V_{ij}}}=\frac{S}{\sigma_{ij}},$
and the bound (3.9) follows. ∎
Corollary 3.10 (Asymptotic normality).
Under the conditions of Theorem 3.9,
$\frac{U_{ij}^{(-j)}}{\sqrt{V_{ij}}}\;\overset{d}{\longrightarrow}\;\mathcal{N}(0,1)\qquad\text{uniformly over }1\leq i,j\leq n.$
Proof.
Immediate from Theorem 3.9, since $h_{n}^{-1/2}\to 0$. ∎